OpenStudy (anonymous):

How would I write a vector equation when given parametric equations? For example, using the parametric equations x=-2-t and y=4+2t, teR

OpenStudy (loser66):

isolate t in both x and y let them equal for example, x = -2-t --> t = -2 -x do the same with y let them equal to get the result

OpenStudy (anonymous):

ok so I got t=-2-x and t=-2+y/2. then I had them equal to each other and got -2x=y. what would I do after that?

OpenStudy (loser66):

dat sit. your vector has property that y = -2x, that is the whole line y = -2x. if x =0, y=0 if x =1, y =-2 so far and so on.